# DisCO: Reinforcing Large Reasoning Models with Discriminative Constrained Optimization

### v2

@@ 6.1 Comparison with Baselines
Task Setting. We validate our method on mathematical reasoning tasks. Specifically, we use the DeepScaleR-Preview-Dataset [42] for training, which includes AIME problems from 1984 to 2023, AMC problems before 2023, and questions from the Omni-MATH [21] and Still [45] datasets, totaling approximately 40.3k unique problem-answer pairs. We evaluate models on six benchmark datasets: AIME 2024, AIME 2025, MATH 500 [27, 37], AMC 2023, Minerva [34], and Olympiad Bench (O-Bench) [30]. Following [23, 42], we adopt the pass@1 metric [11] averaged over $k=16$ responses for each question to ensure the reliability of model performances. The metric for each question is calculated as $\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}(o_{i}\ \text{is correct})$, where $o_{i}$ denotes the $i$th generated response. For both the training and evaluation of our method and the baselines (unless otherwise specified), the maximum response length is limited to 8k tokens.
Models. We conduct experiments with fine-tuning two models: DeepSeek-R1-Distill-Qwen-1.5B model and DeepSeek-R1-Distill-Qwen-7B model. Both are distilled reasoning models.
Baselines. We primarily compare our methods with five most recent state-of-the-art reinforcement learning methods, including (1) GRPO [23]; (2) GRPO with an entropy regularization (GRPO-ER) that adds an entropy on probabilities of output tokens as a regularization to prevent entropy collapse, which is used by DeepScalR [42]; (3) Dr. GRPO [40]; (4) DAPO’s objective [79]; (5) TRPA [61]. For a comprehensive evaluation, we also include a set of reasoning models that are trained from the same base model by other studies with various techniques, such as (6) STILL-3-1.5B-preview [12], which adapt GRPO by periodically replacing the reference model after a fixed number of training steps; (7) DeepScaleR(DSR)-1.5B-Preview that uses maximum response length of 24k for training [42]; (8) GRPO-LEAD-7B [82], which extends GRPO by incorporating length-dependent rewards, explicit penalty terms, and difficulty-based advantage reweighting to encourage concise and precise reasoning.
Training Details. For all the methods, we tune the constant learning rate in $[1\mathrm{e}^{-6},2\mathrm{e}^{-6}]$ with AdamW optimizer with weight decay as 0.01. Generally, a learning rate of $2\mathrm{e}^{-6}$ performs better for the 1.5B model, while $1\mathrm{e}^{-6}$ is better for the 7B model. We employ a training batch size of 128, a mini-batch size of 32, and 8 responses for each question. The temperature is set to 0.6 for both training and evaluation, following the usage recommendation from [23]. For GRPO, $\beta$ is set to 0.001 as commonly used [12, 42]. For GRPO-ER, we use a coefficient of 0.001 for the entropy regularization [42]. For DAPO, we set $\epsilon_{low}$ to 0.2 and $\epsilon_{high}$ to 0.28 by following their paper. For our method, $\delta$ is set to $10^{-4}$ based on the empirical observation that the average KL divergence is around $2*10^{-5}$ and $\beta$ is set to $10^{3}$ such that the effective weight of the KL regularization when the constraint is violated by $\delta$ is on the order of $\beta*\delta=0.1$. Since L-ratio and log-L scoring functions have different orders, we choose $\tau=1$ from $\{0.5,1,5\}$ for L-ratio and $\tau=10$ from $\{1,5,10\}$ for log-L scoring function. For fair comparisons, we do not implement Dynamic Sampling [79] for DAPO and other methods, as it introduces approximately three times the sampling cost at each training step. All methods are run for 1,400 steps on 1.5B models and 1,000 steps on 7B models. Evaluations are conducted every 200 steps, and the best performance for each method is reported.
@@
Performance. We evaluate all the models across six mathematics-focused benchmark datasets to demonstrate the effectiveness of DisCO. The results are summarized in Table 3 and  3. From Table 3 for 1.5B models, we can observe our proposed DisCO methods consistently outperform other baselines by a large margin. Notably, DisCO (log-L) with 8k length for both training and inference achieves an 7% average improvement over GRPO and surpasses DeepScaleR-1.5B-Preview that was trained with maximum 24k length and evaluated with 32k length. A similar trend is observed for 7B models (Table 3), where DisCO significantly outperforms all competing approaches.
@@
Training Dynamics. We compare the training dynamics of different methods in terms of training rewards and generation entropy. From Figure 4(b), we can observe that all baselines suffer from premature saturation due to either entropy collapse for GRPO, GRPO-ER, Dr. GRPO or excessive entropy growth of DAPO, which leads to an early deterministic or highly random policy. The entropy collapse phenomenon is also observed by [79, 26, 42]. TRPA that uses a KL divergence regularization is also observed with instability in the generation entropy in later steps (around 1100 for 1.5B model and around 800 for 7B model). In contrast, our methods with the two scoring functions are most stable, with training rewards kept increasing and generation entropy maintained around 0.25.
@@
Figure 4: Training dynamics of different methods: left two are for fine-tuning 1.5B model and right two are for fine-tuning 7B model. (a), (c) plot the training reward (averaged over generated outputs for questions used in each step) vs the number of training steps (cf. Algorithm 1); (b), (d) plot the generation entropy vs training steps.
@@ 6.2 Ablation Studies
DisCO vs DisCO-b. Figure LABEL:fig:score (left) compares DisCO with DisCO-b using the L-ratio scoring function for training 7B models when evaluated at 1000 steps. The comparison clearly demonstrates the significant improvements of DisCO over DisCO-b, especially on AIME datasets. We also compare DisCO with DisCO-b for other settings in Appendix A.1, and observe that DisCO is consistently better than DisCO-b on average in all settings. It is also notable that DisCO-b with different scoring functions are also better than other baselines trained or evaluated by us in Table 3 and Table 3.
Clipping vs Non-Clipping scoring functions. We compare non-clipping scoring functions L-ratio, log-L with clipped L-ratio (5) in our DisCO-b approach for training 1.5B models in Figure LABEL:fig:score (middle two). We can see that clipped L-ratio causes the entropy collapse yielding worse performance than non-clipping scoring functions.
Sensitivity of hyperparameter $\tau$. We investigate the sensitivity of DisCO to hyperparameter $\tau$ on training 1.5B models. We run DisCO for 1000 steps with different $\tau\in\{0.5,1,5\}$ for L-ratio and different $\tau\in\{1,5,10\}$ for log-L. The result shown in Figure LABEL:fig:score (right) indicate that DisCO is not sensitive to $\tau$ in these ranges.
@@ 7 Conclusion
In this work, we have proposed a novel discriminative constrained optimization framework for reinforcing large reasoning models motivated by the analysis of the GRPO objective. The proposed framework is grounded in the principle of discriminative learning, avoiding difficulty bias and enhancing training stability with constrained trust region optimization. The experiments on mathematical reasoning demonstrated the significant superiority of our approaches, compared with GRPO and its recent variants.
@@ A.1 Detailed comparison between DisCO and DisCO-b
In this part, we compare DisCO and DisCO-b with different score functions on different models. As shown in Figure LABEL:fig:disco_com, DisCO consistently demonstrates better performance compared to DisCO-b across all settings, with higher average scores observed in each case. This consistent advantage highlights the effectiveness of the full DisCO framework. Additionally, it is worth emphasizing that even the DisCO-b variants, with L-ratio or log-L scoring functions, outperform all other baseline methods that are presented in Table 3 and Table 3. These results collectively underscore the robustness and general effectiveness of the DisCO approach.
@@ B.1 Proof of Proposition 1
8:        Compute KL divergence estimator by $\hat{\mathbb{D}}_{KL}=\frac{1}{\sum_{q\in\mathcal{B}_{m}}\sum_{o\in S_{q}}|o|}\sum\limits_{q\in\mathcal{B}_{m}}\sum\limits_{o\in S_{q}}\sum\limits_{t=1}^{|o|}\log\frac{\pi_{\theta_{old}}(o_{t}|q,o_{<t})}{\pi_{\theta}(o_{t}|q,o_{<t})}$
@@ B.2 Connection between discriminative objectives and surrogate objectives in RL
The score function L-ratio is inspired by the same principle as the surrogate objective in TRPO [56]. TRPO aims to maximize the following objective subject to a constraint:
@@
In the reinforcement learning literature, vanilla policy gradient methods also gain significant attention due to their simplicity and remarkable performance. The vanilla policy gradient methods work by computing an estimator of the policy gradient and plugging it into a stochastic gradient algorithm. The most commonly used surrogate objective function for gradient estimator has the form:
@@
Table 4: Weighted discriminative objectives and their scoring functions $s^{+}(o,q)$ and $s^{-}(o,q)$ for different methods.
@@
TRPA $\omega(q)=1,\quad\ell(s)=\log\sigma(\beta(o)s)$

### v4

@@ 6.1 Comparison with Baselines
Performance. We evaluate all the models across six mathematics-focused benchmark datasets to demonstrate the effectiveness of DisCO. The results are summarized in Table 4,  4 and 4. From Table 4 for Q1.5B models, we can observe that our proposed DisCO methods consistently outperform other baselines by a large margin. Notably, DisCO (log-L) with 8k length for both training and inference achieves an 7% average improvement over GRPO and surpasses DeepScaleR-1.5B-Preview that was trained with maximum 24k length and evaluated with 32k length. A similar trend is observed for Q7B models and L8B models (Table 4 and Table 4), where DisCO significantly outperforms all competing approaches.
@@
Training Dynamics. We compare the training dynamics of different methods in terms of training rewards and generation entropy. From Figure 4 for fine-tuning Qwen-1.5B and Qwen-7B models, we can see that all baselines suffer from premature saturation due to either entropy collapse for GRPO, GRPO-ER, Dr. GRPO or excessive entropy growth of DAPO, which leads to an early deterministic or highly random policy. The entropy collapse phenomenon is also observed by yu2025dapo; skywork-or1-2025; deepscaler2025. TRPA that uses a KL divergence regularization is also observed with instability in the generation entropy in later steps (around 1100 for Q1.5B model and around 800 for Q7B model). In contrast, our methods with the two scoring functions are most stable, with training rewards kept increasing and generation entropy maintained around 0.22. We also include training dynamics for the L8B model in Appendix A.2, which follows a similar trend.
@@
Figure 4: Training dynamics of different methods: left two are for fine-tuning DeepSeek-R1-Distill-Qwen-1.5B model and right two are for fine-tuning DeepSeek-R1-Distill-Qwen-7B model. (a), (c) plot the training reward (averaged over generated outputs for questions used in each step) vs the number of training steps (cf. Algorithm 1); (b), (d) plot the generation entropy vs training steps.
(a) Ablation studies: left for comparing DisCO vs DisCO-b; middle and right for comparing clipping with non-clipping scoring functions.
@@ 6.2 Ablation Studies
DisCO vs DisCO-b. Figure 7(a) (left) compares DisCO with DisCO-b using the L-ratio scoring function for training Q7B models for 1000 steps. The comparison clearly demonstrates the significant improvements of DisCO over DisCO-b, especially on the difficult AIME datasets. We also compare DisCO with DisCO-b for other settings in Appendix A.1, and observe that DisCO is consistently better than DisCO-b on average in all settings. It is also notable that DisCO-b with different scoring functions are also better than other baselines trained or evaluated by us in Table 4 and Table 4.
Clipping vs Non-Clipping scoring functions. We compare non-clipping scoring functions L-ratio, log-L with clipped L-ratio (5) in our DisCO-b approach for training Q1.5B models in Figure 7(a) (middle and right). For the clipped L-ratio, we adopt two versions: one with $\epsilon_{high}=0.2$ to align with GRPO objective, and another with $\epsilon_{high}=0.28$ similar to DAPO objective. We can see that clipped L-ratio with $\epsilon_{high}=0.2$ causes the entropy collapse while clipped L-ratio with $\epsilon_{high}=0.28$ leads to excessively high entropy level, both yielding worse performance than non-clipping scoring functions.
KL Regularization vs Constrained Optimization. We investigate the advantages of constrained optimization over KL regularization for DisCO. Specifically, the KL regularization weight is set to the commonly used 0.001 chen2025empirical; deepscaler2025. As shown in Figure 7(b) (left), constrained optimization performs better than KL regularization on both Q1.5B and Q7B models. Moreover, during our experiments, we observed that KL regularization leads to instability in training on Q7B models, similar to TRPA, which indicates that KL regularization is not sufficient to stabilize training.
Sensitivity of hyperparameter $\tau$. We study the sensitivity of DisCO to hyperparameter $\tau$ on training Q1.5B models. Similar to above experiments, we run DisCO for 1400 steps with different $\tau\in\{0.5,1,5,10\}$. The result shown in Figure 7(a) (right) indicates that DisCO is not sensitive to $\tau$ in these ranges.
Effect of each design choice. We analyze the individual contribution of each component in DisCO by replacing its components separately with other designs. We experiment on Q1.5B models and compare with (1) DisCO-b that removes hard negative weighting; (2) adding question-level weight bias $\sqrt{p(q)(1-p(q)}$ to DisCO-b, (3) replacing the KL-divergence constraint with a KL-divergence regularization in DisCO-b, and (4) using a clipping scoring function with $\epsilon_{high}=0.2$ in DisCO-b, respectively. From Figure 7(b) (right), we can see that each of our proposed components is important in DisCO’s improvement, where the use of a non-clipping scoring function is of vital importance.
(b) Ablation studies: left for comparing KL regularization vs constrained optimization; middle for sensitivity of DisCO w.r.t. the hyperparameter $\tau$; right for contribution of each component.
@@ 7 Conclusion
In this work, we have proposed a novel discriminative constrained optimization framework for reinforcing large reasoning models, motivated by the analysis of the GRPO objective. The proposed framework is grounded in the principle of discriminative learning, avoiding difficulty bias and enhancing training stability with constrained trust region optimization. The experiments on mathematical reasoning demonstrated the significant superiority of our approaches, compared with GRPO and its recent variants. While this work focuses on binary rewards, future extensions could incorporate discriminative ranking objectives, like listnet, to handle non-binary rewards. It would be interesting to apply the proposed approaches for fine-tuning larger models or other reasoning tasks.
Acknowledgements
We are grateful for the reviewers’ constructive comments. G. Li and T. Yang were partially supported by NSF grant #2306572.
@@ A.1 Detailed comparison between DisCO and DisCO-b
In this part, we compare DisCO and DisCO-b with different score functions on different models. As shown in Figure 7(c), DisCO consistently demonstrates better performance compared to DisCO-b across all settings, with higher average scores observed in each case. This consistent advantage highlights the effectiveness of the full DisCO framework. Additionally, it is worth emphasizing that even the DisCO-b variants, with L-ratio or log-L scoring functions, outperform all other baseline methods that are presented in Table 4 and Table 4. These results collectively underscore the robustness and general effectiveness of the DisCO approach.
(c) Comparison between DisCO-b and DisCO on different models with different score functions.
A.2 Training dynamics for fine-tuning 8B model.
In this part, we present the training dynamics of different methods for fine-tuning the DeepSeek-R1-Distill-Llama-8B model in Figure 7. Similar to observation in Figure 4 for fine-tuning 1.5B and 7B models, we can see that GRPO, GRPO-ER, and Dr. GRPO still suffer from entropy collapse while DAPO leads to excessive entropy growth, all accompanied by premature saturation in training reward. TRPA with a KL divergence regularization is also observed with instability in the training, indicating the insufficiency of KL regularization to stabilize training. In contrast, our methods with the two scoring functions and the KL constraint demonstrate the greatest stability, with training rewards continuing to rise and generation entropy remaining around 0.2.
Figure 7: Training dynamics of different methods for fine-tuning DeepSeek-R1-Distill-Llama-8B model. (a) plots the training reward (averaged over generated outputs for questions used in each step) vs the number of training steps; (b) plots the generation entropy vs training steps.
A.3 Experiments on DAPO-Math-17K dataset.
In order to demonstrate that the improvements achieved by DisCO are fundamental, rather than relying on specific properties of the dataset, we conducted additional experiments on the DAPO-Math-17K dataset [yu2025dapo] using 1.5B models, training them for 1400 steps. As shown in Table 5, DisCO methods still outperform other baselines by a large margin, demonstrating the generalizability of the proposed method to other datasets.
Table 5: Comparison with baseline methods for fine-tuning DeepSeek-R1-Distill-Qwen-1.5B models on DAPO-Math-17K dataset.
[…]
GRPO 8k / 8k 0.342 0.256 0.842 0.672 0.267 0.458 0.473
GRPO-ER 8k / 8k 0.290 0.260 0.852 0.681 0.287 0.463 0.472
Dr. GRPO 8k / 8k 0.300 0.250 0.849 0.705 0.292 0.464 0.477
DAPO 8k / 8k 0.275 0.229 0.812 0.653 0.256 0.441 0.444
TRPA 8k / 8k 0.346 0.279 0.836 0.683 0.281 0.450 0.479
DisCO (L-ratio) 8k / 8k 0.413 0.310 0.874 0.775 0.307 0.495 0.529
DisCO (log-L) 8k / 8k 0.460 0.317 0.873 0.775 0.320 0.502 0.541
@@ B.1 Proof of Proposition 1
8:   Compute KL divergence estimator by $\hat{\mathbb{D}}_{KL}=\frac{1}{\sum_{q\in\mathcal{B}_{m}}\sum_{o\in S_{q}}|o|}\sum\limits_{q\in\mathcal{B}_{m}}\sum\limits_{o\in S_{q}}\sum\limits_{t=1}^{|o|}\log\frac{\pi_{\text{old}}(o_{t}|q,o_{<t})}{\pi_{\theta}(o_{t}|q,o_{<t})}$
@@ B.2 Connection between discriminative objectives and surrogate objectives in RL
The score function L-ratio is inspired by the same principle as the surrogate objective in TRPO [schulman2015trust]. TRPO aims to maximize the following objective subject to a constraint:
@@
In the reinforcement learning literature, vanilla policy gradient methods also gain significant attention due to their simplicity and remarkable performance. The vanilla policy gradient(VPG) methods work by computing an estimator of the policy gradient and plugging it into a stochastic gradient algorithm. The most commonly used surrogate objective function for gradient estimator has the form:
@@
Table 6: Weighted discriminative objectives and their scoring functions $s^{+}(o,q)$ and $s^{-}(o,q)$ for different methods, where $\sigma(\cdot)$ is the sigmoid function.
@@
TRPA $\omega(q)=1,\quad\ell(s)=\log(\sigma(\beta(o)s))$